OpenStudy (anonymous):

The next digit in the series 6, 4, 3, 12/5 will be : a) 7/3 b) 9/4 c) 13/5 d) 2

OpenStudy (campbell_st):

the terms are being multiplied, the pattern is shown below 6 x 2/3 = 4 4 x 3/4 = 3 3 x 4/5 = 12/5 12/5 x 5/6 =

OpenStudy (anonymous):

Thank u . .I got it :)